OpenStudy (anonymous):

solve sqrt 96/xsqrt23=1/2

OpenStudy (anonymous):

\[\sqrt{96}/x \sqrt{23}=1/2\]

OpenStudy (anonymous):

\[\sqrt{\frac{96}{23}}\frac{1}{x} = \frac{1}{2}\] \[\sqrt{\frac{96}{23}} = \frac{x}{2}\] \[2\sqrt{\frac{96}{23}} = x\]

OpenStudy (anonymous):

\[x=2\sqrt{\frac{16*6}{23}}=2*4*\sqrt{\frac{6}{23}}=8\sqrt{\frac{6}{23}}\]

OpenStudy (anonymous):

I continued from where Ishaan left off. Using 96=16*6

OpenStudy (anonymous):

http://www.twiddla.com/606133

OpenStudy (anonymous):

and since 16=4^2 , It can be taken out of the radical and become a 4.

OpenStudy (anonymous):

\[\frac{\sqrt{96}}{x\sqrt{23}} = \frac{1}{2}\]Multiply both sides by 2x..\[\frac{\sqrt{96}}{\cancel{x}\sqrt{23}} \cdot 2\cancel{x} = \frac{1}{\cancel{2}}\cdot \cancel{2}x\]\[\implies \frac{2\sqrt{96}}{\sqrt{23}} = x\]\[\implies x = \frac{2\cdot \sqrt{16 \cdot 6}}{\sqrt{23}} = \frac{2\cdot 4 \cdot \sqrt{6}}{\sqrt{23}} = 8\sqrt{\frac{6}{23}}\]